OpenStudy (pottersheep):

Is the only way to solve these kinds of questions to graph it and see visually? (Grade 11 beginner)

OpenStudy (pottersheep):

For exp. graph y = x^2, y=(2x)^2 , y = (1/2x)^2 find invariant points

OpenStudy (pottersheep):

would I have to graph those to find the invariant points?

OpenStudy (anonymous):

do you know the definition of an invariant point?

OpenStudy (mathteacher1729):

How are "invariant points" defined?

OpenStudy (pottersheep):

Points that are the same?

OpenStudy (anonymous):

the same as what? Look it up.

OpenStudy (mathteacher1729):

So basically you have y = (kx)^2 for k = (some range of numbers) and you want to find the values of x which always yield the same values of y, regardless of the values of k? If yes, the only point is (0,0) for this graph. (see attached).

OpenStudy (mathteacher1729):

made with: Geogebra --> http://www.geogebra.org

OpenStudy (pottersheep):

a fixed point = invariant point